OpenStudy (anonymous):

Find the values of x that satisfies the following inequality: 4x+3/x^3+2x^2+3x+6<2/x-3

OpenStudy (anonymous):

\[\frac{ 4x+3 }{ x^3+2x^2+3x+6 } <\frac{ 2 }{ x-3 }\]

OpenStudy (anonymous):

hw doI begin?

OpenStudy (anonymous):

carry out cross multiplication i.e. (4x+3)(x-3)<2(x^3+2x^2+3x+6)

OpenStudy (anonymous):

ok then...

OpenStudy (anonymous):

cancel the like terms from both sides having same sign and solve the remaining problem

OpenStudy (anonymous):

\[\frac{4x+3}{x^3+2x^2+3x+6}< \frac{2}{x-3}\] i.e. \[(4x+3)(x-3)<2(x^3+2x^2+3x+6)\] \[4x(x-3)+3(x-3)<2x^3+4x^2+6x+12\] \[\rightarrow 4x^2-12x+3x-9<2x^3+4x^2+6x+12\] in the next step we will transfer variables on LHS and constants on RHS \[4x^2-12x+3x-2x^3-4x^2-6x<+12 +9\] \[\rightarrow -15x-2x^3<21 \rightarrow 2x^3+15x<-21 \rightarrow 2x^3+15x+21<0\]

OpenStudy (anonymous):

I am not able to see the fraction in proper order. What shld I do?

OpenStudy (anonymous):

[\frac{ -15x-2^3-21 }/{ (x+2)(x^2+3)(x+3)} < 0\]\]

OpenStudy (asnaseer):

where are you stuck?

OpenStudy (anonymous):

nw, hw to solve for x values? is it x = -3,-2,3 not be included

OpenStudy (asnaseer):

the inequality you should end up with is:\[2x^3+15x+21<0\]solving this depends on what you have been taught in class regarding cubic equations.

OpenStudy (anonymous):

is it use factor theorem?

OpenStudy (anonymous):

I need help...

OpenStudy (asnaseer):

use whatever theorem(s) you have been taught in class.